OpenStudy (andrea):

a chord of a circle is 12 mm long and is 3 mm from the center of the circle. what is the area of the circle?

OpenStudy (anonymous):

you can use pythagorus to find the radius of the circle: 

OpenStudy (anonymous):

then we know area is given by the formula \[A = \pi r^2\]

OpenStudy (anonymous):

If the chord is 12 mm, then 1/2 of a chord is 6 mm. Then, to find the radius, you do r^2 = (3^2) + 6^2 = 45. Area = pi*r^2, so A = 45pi